A fan blade rotates with angular velocity given by ωz(t)= γ − β t2, where γ = 5.00 rad/s and β = 0.800 rad/s3.

a- Calculate the angular acceleration as a function of time.

b- Calculate the instantaneous angular acceleration αz at t = 3.00 s and the average angular acceleration aav-z for the time interval t=0 to t0 = 3.00 s.

How do these two quantities compare? If they are different, why are they different?



1
Expert's answer
2022-01-14T09:48:10-0500

a) The angular acceleration is the first derivative of velocity by time:


"\alpha (t)=0-2\beta t=-1.6t."


b) The instantaneous value of acceleration:


"\alpha (3)=-4.8\text{ rad/s}^2."


The average angular acceleration is the sum of angular accelerations divided by time it took:


"\alpha_\text{avg}=\dfrac{\alpha (3)+\alpha (0)}{t}=-1.6\text{ rad/s}^2."


The values are different because the average value is averaged along a certain time, while the instantaneous value is the value at a certain moment.